OpenStudy (okagurao):

Which value is a solution to the open sentence? 6x – 3 = 39 A. 4 B. 5 C. 6 D. 7

OpenStudy (okagurao):

i will fan whoever helps

OpenStudy (anonymous):

6x-3=39 6x=42 x=42/6 x=7